# SAGA: Synthetic Audit Log Generation for APT Campaigns

### v1

@@ I Introduction
However, labeling large volumes of audit logs over long periods poses challenges, especially when APT attacks lack obvious indicators. Researchers often rely on publicly available datasets such as the DARPA Transparent Computing Engagement [8], Operationally Transparent Cyber (OpTC)[9], and LANL’s dataset[10], as well as smaller datasets that are rarely shared due to privacy and proprietary concerns. This shortage of labeled logs limits the comprehensive testing of cybersecurity solutions and restricts the application of AI/ML in advancing cybersecurity studies [11, 12].
@@ II-C Running Example
We refer to APT28 threat intelligence reports [16] to construct the attack scenario, illustrated in Figure 2, which serves as the running example in this work. Figure 2 a outlines its attack lifecycle, while Figure 2 b shows the corresponding APT audit log. In the initial stages of the attack, phishing emails containing malicious attachments are sent to the victim to download the attachment IOC_09_11.rar. This activity is designated as the Technique T1566.001 Spearphishing Attachment, as (1) in Figure 2 b, of the attack lifecycle stage Initial Compromise. This malicious attachment exploits a vulnerability in WinRAR versions below 6.23 (as mentioned in CVE-2023-38831). When the victim clicks on the IOC_09_11.rar file, it triggers subsequent malicious activities designated as T1204.002 User Execution: Malicious File of the attack lifecycle Establish Foothold, as (2) in Figure 2 b.
@@ III SAGA Generative Model
The SAGA generative model involves collecting audit logs from a red-team emulator to create a repository of attack pattern templates with fine-grained labels. These templates are then used, following lifecycle selection instructions, to generate synthetic audit logs that embed APT campaigns. Figure 3 provides an overview and the workflow of SAGA, which will be discussed in detail in the following subsections.
@@ III-A Attack Pattern Collection and Labeling
The first task of SAGA is to collect labeled audit logs, as depicted in Figure 3. The Red Team Platform, such as CALDERA [17, 18] and Atomic Red Team [19], simulates real-world adversaries in a controlled environment. In this study, we use CALDERA to emulate real-world attacks by adversaries. Our choice is based on its open source nature and alignment with the MITRE ATT&CK framework. To deploy CALDERA, we configure a master server and set up Remote Access Tool (RAT) clients on a compromised host. An ability of CALDERA refers to a specific implementation of an ATT&CK Technique, representing a discrete adversarial behavior. After acquiring audit logs from a red team emulator and process monitor, the next step is to extract the relevant attack patterns and assign appropriate Technique labels. CALDERA-generated reports provide details for each executed attack, including process IDs, commands used, and associated technique IDs. By mapping process IDs and commands to the corresponding audit logs, we can accurately assign technique ID to the relevant entries in the logs.
@@ III-B Attack Pattern Templates Abstraction
Each attack pattern corresponds to a specific instance of a Technique with particular parameter values. To enable the generation of a wider variety of attack patterns, it is essential to abstract these patterns into a more abstracted form. These abstractions can then be instantiated with legitimate parameter values, allowing the creation of new attack patterns. This approach facilitates greater flexibility and diversity in the generation of new attack instances.
@@
Figure 5: APT28 Attack pattern template example
@@
Attack Pattern. The entity Attack Pattern is associated with an attack identity and consists of a sequence of events generated by a specific ability, and the categories and descriptors of artifacts presented in the event. In this context, the category provides a broad classification, while the descriptor offers a more specific identification. For example, if an event contains the artifact Firefox, it falls into the category process with the descriptor browser. To support the generalization of attack pattern templates, Table I outlines a collection of categories and descriptors used in SAGA, which can be considered as forming a hypernym relationship. The table outlines six categories, File, Network, Cmdline, Process, Registry, and System, each subdivided into specific descriptors.
Subsequently, additional attack instances can be generated by instantiating these attack pattern templates during the generation of synthetic audit logs. Using the same running example, three system entities (artifacts) are involved in phishing: Firefox accessing Microsoft Exchange and downloading IOC_09_11.rar through the browser. These entities can be substituted with hypernyms (as shown in lines 7-10 of Figure 5), such as replacing Firefox with Process.browser and IOC_09_11.rar with File.phishing.
Prerequisite. The entity Prerequisite specifies conditions related to the system resources required for the execution of an attack pattern, such as process.browser and file.phishing (as shown in lines 12-15 of Figure 5) of the attack pattern template T1566.001 of APT28. To generate a specific malicious audit log (i.e., an attack pattern), a campaign information table records previously executed attacks and the system entities utilized, as well as the system entities provided by the external environment, to facilitate Prerequisite matching. An attack pattern is generated only when all Prerequisite are satisfied; more details are given in Section III-C2.
Outcome. The entity Outcome lists the artifacts produced after executing this attack, such as a phishing file (as indicated on line 17 of Figure 5), collected sensitive information, and changes in system state. In a complete APT attack chain, Outcome may serve as prerequisites for subsequent attack patterns. Therefore, defining the Outcome within the template is crucial, as it not only impacts the immediate effects of the attack, but also has long-term implications for future attack strategies and defensive measures.
0:  Attack pattern $A$ and events $E$
0:  Attack pattern template $P$
1:  Let $TAB$ be the table that stores known system entities and their descriptors and categories.
2:  for each event $e$ in $E$ do
3:     for each system entity $s$ in $E$ do
4:        if $s$ already used then
5:           Replace $s$ with its descriptor $d$ from $TAB$.
6:        else
7:           Find its descriptor $d$ and replace $s$.
8:           Set the descriptor as a prerequisite.
19:        end if
10:        Set $d$ as an outcome if it still exists.
11:        Insert $e$ into $P$.
12:     end for
13:  end for
14:  return  $P$
Algorithm 1 Generate Attack Pattern Template
@@ III-C1 Lifecycle Stage Transitions
1. $SA$ :- $C_{m}(SA,MA)$ |||| $C_{b}(SA,BA)$ |||| $MA$ |||| $BA$ |||| $\emptyset$
2. $BA$ :- $EVT$
3. $EVT$ :- $evt+EVT$ |||| $evt$ |||| $\emptyset$
@@ 6. $LC$ :- $IC+EF+(EP^{0+}+IR^{0+}+ML^{0+}+$
8. $MA$ :- $BE+TQ+MA$ |||| $BE+TQ$
9. $BE$ :- $EVT$
10. $TQ$ :- $In(TQ)$ if TQ.prerequisites are satisfied |||| $\emptyset$
@@ III-C2 Attack Template Selection
For each stage of an APT campaign, the number of attacks can be either randomly determined or specified by the user. An attack template (TQ) may also be randomly selected based on the current stage and prerequisites, or as specified by the user. Note that there may be benign events (BE) preceding the attack pattern templates (TQ) that the operation is described in line 8 of Table II. In line 8, malicious audit logs (MA) consist of either a sequence of background events (BE), the selected attack pattern instance (TQ), malicious audit logs (MA), or sequences of background events (BE) and the selected attack pattern instance (TQ). Next, the abstracted arguments of the template are replaced with artifacts to generate diverse audit events, as the instantiation of an attack pattern described in line 10 ($In(TQ)$) and elaborated in Subsection III-C3. The construction of the malicious audit log continues until all stages have been completed.
A campaign information table, as depicted in Figure 4, is created to maintain the information about the executed attack templates, such as their associated lifecycle stage, as well as the used system entities and their statuses. Given the current stage, the prerequisites of a template are examined against the campaign information table to determine whether they are satisfied. For example, if a threat actor uses a phishing email template during the initial compromise stage, the campaign information table will record the information of the phishing email template, including the entity File.phishing. In the second stage (EF), the templates for this stage will refer to the campaign information table to check if their prerequisites are met. Only if the prerequisites are met can they be considered potential templates for this APT campaign.
0:     Lifecyle $\{LC\}$, Attack pattern templates $\{TQ\}$, Benign audit log $\{BA\}$, Background events $\{BG\}$
10:     Synthetic audit events $SA$
1:  while while $LC_{i}$ in $LC$ exists do
2:     for each $stage$ in $LC_{i}$ do
3:        // Include an attack pattern template $TQ$ in the $stage$. Line 8 of Table II.
4:        $MA$ = $BE+TQ+MA$ |||| $BE+TQ$
25:        where $TQ$ is an instantiated template, as Line 10 of Table II. $BE$ is a short list of benign events.
@@
9:  $SA$ = $BG$
@@
14:     end for // Insert event $evt$ to synthetic audit log.
15:     $SA$ = $SA\cup BA_{i}$
@@
19:     for each $TQ_{j}$ in $\{MA_{i}\}$ do
20:        for each $evt$ in $\{TQ_{j}\}$ do
21:           $evt.start={TQ_{j-1}}.end+TQ_{j}.lapse+evt.relative$
22:        end for // Insert event $evt$ to synthetic audit log.
23:        $SA$ = $SA\cup TQ_{j}$
24:     end for // Insert technique $TQ$ to synthetic audit log.
425:  end for // Insert malicious audit log $MA$ to synthetic audit log.
@@ III-D Synthetic Audit Log Generation
For the composition of an APT campaign audit log into a synthetic audit log, denoted as $C_{m}(SA,MA)$ in the first line of Table II, it merges each pair of benign events (BE) and attack technique events (TQ) sequentially until all pairs are combined. A $BE$ usually is a short sequence of benign events that occur before the particular Technique (TQ). The composition of $BE$ follows the same method as that for $BA$. The composition of $TQ$, which refers to the audit log events representing a particular Technique, provides flexibility in both the start time and duration. This means that a Technique audit log can begin at any time and span any length. The start time of a Technique audit log is set to the end time of the preceding Technique plus the lapse time $TQ.lapse$, which can be specified by the user or decided randomly by SAGA. The procedure is described in lines 15 to 22 of Algorithm 2. Figure 6 depicts a composition of 3 benign audit logs and 4 APT audit logs. Note that the logs may overlap, and the APT audit logs can be extended for any duration.
@@ IV-A Evaluation Settings
Given the limited availability of audit logs containing labeled with Techniques, we used the attack pattern templates to create APT campaigns. These scenarios include eight APT attack campaigns (“known” group) based on cyber threat intelligence reports from real-world APT campaigns, including Higaisa [25], admin338 [26], APT28 [16], FIN7 [27], CobaltGroup [28], Gamaredon [29], Patchwork [30], and GorgonGroup [31]. Moreover, using SAGA random generation capability, we produced 20 APT campaigns (“random generated group”). From these 28 APT campaigns, we randomly selected 3 to create new composite scenarios involving multiple APT campaigns targeting a single victim host, resulting in 10 composite campaigns (“composite” group). To demonstrate the versatility and flexibility of SAGA, each of the above APT campaigns was generated over three different time periods: 15 minutes, 1 hour, and 1 day. Labeling. For each generated APT campaign, SAGA labels malicious events with APT stages, Techniques, abilities, manipulated system entities, and other operation-related attributes. In addition, it includes threat actors, when applicable. Malicious audit events are labeled using the BIO2 scheme [32] covering 7 attack stages, 80 Techniques, and 169 abilities, as shown in Table S1 of Supplementary Material A.
@@
Evaluation Metrics. We adopt three evaluation metrics for different tasks: APT intrusion detection, Technique hunting, and APT campaign attribution. The APT intrusion detection task aims to identify any APT intrusion or anomalies within the dataset under examination. Technique hunting seeks to identify Techniques violated, based on MITRE ATT&CK definitions. APT campaign attribution aims to determine the most probable APT campaign present in the dataset. To evaluate performance, we use standard classification metrics: precision (P), recall (R), F1 score (F1), and accuracy (ACC).
@@
For APT intrusion detection, we frame it as a binary classification problem, determining whether a malicious event exists in the audit log. Since different target systems or methods apply their own tests to detect APT intrusion on an event, a graph, or a time window interval, we adapt to their respective evaluation design to assess performance. Technique hunting is treated as a multi-class classification task, where each classifier determines whether a specific Technique has been employed in one or more events. For APT campaign attribution, we analyze the Techniques identified from the Technique hunting task to determine the most likely APT campaign for the audit log. We rank the ground truth against the predictions and employ Top-K measurement to evaluate the performance of the APT campaign attribution.
@@
Established Baselines For APT intrusion detection, we demonstrate the usefulness of generated attack datasets using three types of systems: event-based (AirTag [6]), graph-based (Unicorn [33]), and time window-based (KAIROS [7]) intrusion detection systems. The reasons for choosing the three systems are not only because they are established baselines, but also because their codes are publicly available and the systems can be installed and executed without issues.
@@
For Technique hunting, we used the Sigma rules [34], an open signature-based rulebase contributed by numerous security practitioners, and SFM [35], a deep learning approach that provides Technique hunting functionality. Sigma rules help bridge the gap caused by the lack of a standardized benchmark to represent fine-grained attack patterns, enabling the identification of relevant log events. A major portion of these rules align with the MITRE ATT&CK framework with 70 rules mapped to the Techniques identified in this study. SFM integrates three key technologies: anomaly detection to mitigate the impact of data imbalance, a multi-class classifier for Technique hunting, and subgraph matching for APT campaign attribution. SFM was selected because it was one of the few systems specifically designed to address both the Technique hunting task and APT campaign attribution, making it particularly suited for the evaluation of complex attack scenarios. All models and methods were installed with configurations as close as possible to those described in their original publications. Code modifications were only made to ensure compatibility with the structure of the synthetic audit logs.

### v2

@@ Abstract
publicationid: pubid: © 2025 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
However, labeling large volumes of audit logs over long periods poses challenges, especially when APT attacks lack obvious indicators. Researchers often rely on publicly available datasets such as the DARPA Transparent Computing Engagement [8], Operationally Transparent Cyber (OpTC)[9], and LANL’s dataset[10], as well as smaller datasets that are rarely shared due to privacy and proprietary concerns. Moreover, some datasets provide fine-grained labels on malicious entities [5] or processes [11], while others [8, 9] only report attack scenarios over time intervals, without linking them to specific system events. This shortage of labeled logs limits a deeper understanding of attack behaviors and restricts the development of AI/ML-based cybersecurity defense studies [12, 13].
@@ II-C Running Example
We refer to APT28 threat intelligence reports [17] to construct the attack scenario illustrated in Figure 2, which serves as the running example in this work. Figure 2 a outlines its attack lifecycle, while Figure 2 b shows the corresponding APT audit log. In the initial stages of the attack, phishing emails containing malicious attachments are sent to the victim to download the attachment IOC_09_11.rar. This activity is designated as the Technique T1566.001 Spearphishing Attachment, as (1) in Figure 2 b, of the attack lifecycle stage Initial Compromise. This malicious attachment exploits a vulnerability in WinRAR versions below 6.23 (as mentioned in CVE-2023-38831). When the victim clicks on the IOC_09_11.rar file, it triggers subsequent malicious activities designated as T1204.002 User Execution: Malicious File of the attack lifecycle Establish Foothold, as (2) in Figure 2 b.
@@ III SAGA Generative Model
The SAGA generative model involves collecting audit logs from a red-team emulator to create a repository of attack pattern templates with fine-grained labels. These templates are then used, following lifecycle selection instructions, to generate synthetic audit logs that embed APT campaigns. Figure 3 provides an overview and the workflow of SAGA, which will be discussed in detail in the following subsections. To mitigate the risk of detection models overfitting on artificial patterns, SAGA introduces variability in both attack implementation in Subsection III-A and artifact representation in Subsection III-C3. In Subsection III-A, SAGA collects attack pattern from a red-team emulation platform that supports multiple abilities per Technique, enabling the generation of diverse execution paths and behavioral variants. In Subsection III-C3, SAGA incorporates entity-level variability by replacing generalized entities with hyponyms derived from real-world malware samples and synthetic variants. This combination of implementation diversity and entity-level variation helps generate a wide range of realistic log instances.
@@ III-A Attack Pattern Collection and Labeling
The first task of SAGA is to collect labeled audit logs, as depicted in Figure 3. The Red Team Platform, such as CALDERA [18, 19] and Atomic Red Team [20], simulates real-world adversaries in a controlled environment. In this study, we use CALDERA to emulate real-world attacks by adversaries. Our choice is based on its open source nature and alignment with the MITRE ATT&CK framework. To deploy CALDERA, we configure a master server and set up Remote Access Tool (RAT) clients on a compromised host. An ability of CALDERA refers to a specific implementation of an ATT&CK Technique, representing a discrete adversarial behavior. After acquiring audit logs from a red team emulator and process monitor, the next step is to extract the relevant attack patterns and assign appropriate Technique labels. CALDERA-generated reports provide details for each executed attack, including process IDs, commands used, and associated technique IDs. We employ a semi-automated labeling method to annotate audit logs based on handcrafted rules tailored to specific attack techniques.
@@
When given a CALDERA-generated report, SAGA uses the attack-related Process ID (PID) as the root to construct a comprehensive process family tree that captures all associated processes. By mapping these process IDs to the corresponding audit logs and applying handcrafted rules, SAGA can accurately assign the appropriate technique ID to the relevant log entries.
@@ III-B Attack Pattern Templates Abstraction
Each attack pattern corresponds to a specific instance of a Technique with particular parameter values. To enable the generation of a wider variety of attack patterns, it is essential to abstract these patterns into a more abstract form. These abstractions can then be instantiated with legitimate parameter values, allowing for the creation of new attack patterns. This approach facilitates greater flexibility and diversity in the generation of new attack instances.
@@
Figure 5: APT28 Attack pattern template example
0: Attack pattern AA and events $E$
0: Attack pattern template $P$
1: Let TABTAB that stores known system entities and their descriptors and categories, as defined in Table 1.
2: Initialize $P$ with the stage and Technique, as described in Subsection ”Attack Identification”.
3:for each event $e$ in $E$ do
4:  for each system entity $s$ in $E$ do
5:   if $s$ already used then
6:    Replace $s$ with its descriptor $d$ from TABTAB.
7:   else
8:    Find its descriptor $d$ and replace $s$.
9:    Set the descriptor as a prerequisite.
10:   end if
11:   Set $d$ as an outcome if it still exists.
12:   Insert $e$ into $P$.
13:  end for
14:end for
15:return $P$
Algorithm 1 Generate Attack Pattern Template
@@
Attack Pattern. The entity Attack Pattern is associated with an attack identity and consists of a sequence of events generated by a specific ability, categories and descriptors of artifacts presented in the event. In this context, the category provides a broad classification, while the descriptor offers a more specific identification. For example, if an event contains the artifact Firefox, it falls into the category process with the descriptor browser. To support the generalization of attack pattern templates, Table I outlines a collection of categories and descriptors used in SAGA, which can be considered as forming a hypernym relationship. The table outlines six categories, File, Network, Cmdline, Process, Registry, and System, each subdivided into specific descriptors. Refer to lines 3-6 of Algorithm 1.
Subsequently, additional attack instances can be generated by instantiating these attack pattern templates during the generation of synthetic audit logs. Using the same running example, three system entities (artifacts) are involved in phishing: Firefox accessing Microsoft Exchange and downloading IOC_09_11.rar through the browser. These entities can be substituted with hypernyms (as shown in line 12 of Algorithm 1 and lines 7-10 of Figure 5), such as replacing Firefox with Process.browser and IOC_09_11.rar with File.phishing.
Prerequisite. The entity Prerequisite specifies conditions related to the system resources required for the execution of an attack pattern(lines 8-9 of Algorithm 1), such as process.browser and file.phishing (as shown in lines 12-15 of Figure 5) of the attack pattern template T1566.001 of APT28. To generate a specific malicious audit log (i.e., an attack pattern), a campaign information table records previously executed attacks and the system entities utilized, as well as the system entities provided by the external environment, to facilitate Prerequisite matching. An attack pattern is generated only when all Prerequisite are satisfied; more details are given in Section III-C2.
Outcome. The entity Outcome lists the artifacts produced after executing this attack (line 11 of Algorithm 1), such as a phishing file (as indicated on line 17 of Figure 5), collected sensitive information, and changes in system state. In a complete APT attack chain, Outcome may serve as prerequisites for subsequent attack patterns. Therefore, defining the Outcome within the template is crucial, as it not only impacts the immediate effects of the attack, but also has long-term implications for future attack strategies and defensive measures.
@@ III-C1 Lifecycle Stage Transitions
1. SASA :- $C_{m}(SA,MA)$ $|$ $C_{b}(SA,BA)$ $|$ MAMA $|$ BABA $|$ $\emptyset$
2. BABA :- EVTEVT
3. EVTEVT :- $evt+EVT$ $|$ evtevt $|$ $\emptyset$
@@ 6. LCLC :- $IC+EF+(EP^{0+}+IR^{0+}+ML^{0+}+$
8. MAMA :- $BE+In(TQ)+MA$ $|$ $BE+In(TQ)$
9. BEBE :- EVTEVT
10. $In(TQ)$ :- $In(TQ)$ if TQ.prerequisites are satisfied $|$ $\emptyset$
  • Notation: $SA$ = synthetic audit log, $MA$ = malicious audit log, $BA$ = benign audit log, $LC$ = APT lifecycle, $IC$ = initial compromise, $EF$ = establish foothold, $EP$ = escalate privileges, $IR$ = internal reconnaissance, $ML$ = move laterally, $MP$ = maintain presence, $CM$ = complete mission, $BE$ = a sequence of benign events, $EVT$ = event sequence, $TQ$ = attack template, $In(TQ)$ = attack pattern instance
@@ III-C2 Attack Template Selection
For each stage of an APT campaign, the number of attacks can be either randomly determined or specified by the user. An attack template (TQ) may also be randomly selected based on the current stage and prerequisites, or as specified by the user. Note that there may be benign events (BE) preceding the attack pattern templates (TQ) that the operation is described in line 8 of Table II. In line 8, malicious audit logs (MA) consist of either a sequence of benign events (BE), the selected attack pattern instance ($In(TQ)$), malicious audit logs (MA), or sequences of benign events (BE) and the selected attack pattern instance ($In(TQ)$). Next, the abstracted arguments of the template are replaced with artifacts to generate diverse audit events, as the instantiation of an attack pattern described in line 10 ($In(TQ)$) and elaborated in Subsection III-C3. The construction of the malicious audit log continues until all stages have been completed.
A campaign information table, as depicted in Figure 4, is created to maintain information about the executed attack templates, including their associated lifecycle stage, as well as the system entities used and their statuses. Given the current stage, the prerequisites of a template are examined against the campaign information table to determine whether they are satisfied. For example, if a threat actor uses a phishing email template during the initial compromise stage, the campaign information table will record the details of the phishing email template, including the entity File.phishing. In the second stage (EF), the templates for this stage will refer to the campaign information table to check whether their prerequisites are met. Only if the prerequisites are met can they be considered potential templates for this APT campaign.
0:    Lifecyle $\{LC\}$, Attack pattern instances $\{ln(TQ)\}$, Benign audit log $\{BA\}$
10:    Synthetic audit events SASA
1:while while $LC_{i}$ in LCLC exists do
2:  for each stagestage in $LC_{i}$ do
3:   // Include an attack pattern instance $ln(TQ)$ in the stagestage. Line 8 of Table II.
4:   MAMA = $BE+In(TQ)+MA$ $|$ $BE+In(TQ)$
25:   where $In(TQ)$ is an instantiated attack template, as Line 10 of Table II. BEBE is a short list of benign events.
@@
9: Initialize SASA with given background events.
@@
14:  end for // Insert event evtevt to synthetic audit log.
15:  SASA = $SA\cup BA_{i}$
@@
19:  for each $In(TQ_{j})$ in $\{MA_{i}\}$ do
20:   for each evtevt in $\{In(TQ_{j})\}$ do
21:    $evt.start=In(TQ_{j-1}).end+In(TQ_{j}).lapse+evt.relative$
22:   end for // Insert event evtevt to synthetic audit log.
23:   SASA = $SA\cup In(TQ_{j})$
24:  end for // Insert technique $In(TQ)$ to synthetic log.
425:end for // Insert malicious audit log MAMA to synthetic audit log.
@@ III-D Synthetic Audit Log Generation
For the composition of an APT campaign audit log into a synthetic audit log, denoted as $C_{m}(SA,MA)$ in the first line of Table II, it merges each pair of benign events (BE) and attack technique events ($ln(TQ)$) sequentially until all pairs are combined. A $BE$ is usually a short sequence of benign events that occurs before the particular Technique ($ln(TQ)$). The composition of $BE$ follows the same method as that for $BA$. The composition of $TQ$, which refers to the audit log events representing a particular Technique, provides flexibility in both the start time and duration. This means that a Technique audit log can begin at any time and span any length. The start time of a Technique audit log is set to the end time of the preceding Technique plus the lapse time $ln(TQ).lapse$, which can be specified by the user or determined randomly by SAGA. The procedure is described in lines 15 to 22 of Algorithm 2.
Figure 6 depicts a composition of 3 benign audit logs and 4 APT audit logs. Note that the logs may overlap and the APT audit logs can be extended for any duration. SAGA explicitly leverages process identifiers, parent-child process relationships, and event timestamps to preserve the underlying process structure, temporal ordering, and causal dependencies that are critical for contextual consistency in audit logs. To avoid unrealistic combinations of events, SAGA maintains semantic coherence at both the process and artifact levels. For example, when two events involve shared entities, such as a file, registry key, or named pipe, SAGA preserves these dependencies to ensure that the resulting logs reflect plausible and continuous behavior. As illustrated in Subsection II-C, a user downloads a malicious attachment IOC_09_11.rar via a browser (T1566.001: Spearphishing Attachment), and subsequently opens it using Microsoft Word (T1204.002: Malicious File Execution). Although these steps span separate processes, their interaction through a shared artifact forms a coherent and realistic attack sequence. In this way, SAGA ensures that injected attack behaviors are not only chronologically plausible but also contextually consistent with normal system operations, mitigating the risk of generating disjointed or unrealistic event flows.
@@ IV-A Evaluation Settings
Given the limited availability of audit logs labeled with Techniques, we used the attack pattern templates to create APT campaigns. These scenarios include eight APT attack campaigns (“known” group) based on cyber threat intelligence reports from real-world APT campaigns, including Higaisa [26], admin338 [27], APT28 [17], FIN7 [28], CobaltGroup [29], Gamaredon [30], Patchwork [31], and GorgonGroup [32]. Moreover, using SAGA random generation capability, we produced 20 APT campaigns (“random generated group”). From these 28 APT campaigns, we randomly selected 3 to create new composite scenarios involving multiple APT campaigns targeting a single victim host, resulting in 10 composite campaigns (“composite” group). To demonstrate the versatility and flexibility of SAGA, each of the above APT campaigns was generated over three different time periods: 15 minutes, 1 hour, and 1 day.
Labeling. For each generated APT campaign, SAGA labels malicious events with APT stages, Techniques, abilities, manipulated system entities, and other operation-related attributes. In addition, it includes threat actors, when applicable. Malicious audit events are labeled using the BIO2 scheme [33] covering 7 attack stages, 80 Techniques, and 169 abilities, as shown in Table S1 of Supplementary Material A.
@@
Evaluation Metrics. We adopt three evaluation metrics for different tasks: APT intrusion detection, Technique hunting, and APT campaign attribution. The APT intrusion detection task aims to identify any APT intrusions or anomalies within the dataset under examination. Technique hunting seeks to identify Techniques violated based on MITRE ATT&CK definitions. APT campaign attribution aims to determine the most probable APT campaign present in the dataset. To evaluate performance, we use standard classification metrics: precision (P), recall (R), F1 score (F1), and accuracy (ACC).
@@
For APT intrusion detection, we frame it as a binary classification problem, determining whether a malicious event exists in the audit log. Since different target systems or methods apply their own tests to detect APT intrusion on an event, a graph, or a time window interval, we adapt to their respective evaluation designs to assess performance. Technique hunting is treated as a multi-class classification task, where each classifier determines whether a specific Technique has been employed in one or more events. For APT campaign attribution, we analyze the Techniques identified from the Technique hunting task to determine the most likely APT campaign for the audit log. We rank the ground truth against the predictions and employ a Top-K measurement to evaluate the performance of the APT campaign attribution.
@@
Established Baselines For APT intrusion detection, we demonstrate the usefulness of generated attack datasets using three types of systems: event-based (AirTag [6]), graph-based (Unicorn [34]), and time window-based (KAIROS [7]) intrusion detection systems. The reasons for choosing the three systems are not only that they are established baselines, but also that their codes are publicly available and the systems can be installed and executed without issues.
@@
For Technique hunting, we used the Sigma rules [35], an open signature-based rulebase contributed by numerous security practitioners, and SFM [36], a deep learning approach that provides Technique hunting functionality. We selected Sigma as a baseline because it is a widely adopted, community-driven detection framework, actively maintained by practitioners worldwide. As noted by  [37], Sigma has also been shown to align closely with several commercial rulesets, making it a meaningful and representative point of comparison. A major portion of these rules align with the MITRE ATT&CK framework with 70 rules mapped to the Techniques identified in this study. SFM integrates three key technologies: anomaly detection to mitigate the impact of data imbalance, a multi-class classifier for Technique hunting, and subgraph matching for APT campaign attribution. SFM was selected because it was one of the few systems specifically designed to address both the Technique hunting task and APT campaign attribution, making it particularly suited for the evaluation of complex attack scenarios. All models and methods were installed with configurations as close as possible to those described in their original publications. Code modifications were only made to ensure compatibility with the structure of the synthetic audit logs.